This problem with linear functions is getting me confused, can you help me please? I appreciate it.
In 1991, the life expectancy of males in a certain country was 66.3 years. In 1995, it was 68.9 years. Let E represent the life expectancy in year t and let t represent the number of years since 1991.
The linear function E(t) that fits the data is E(t)=___t + ____ (round to the nearest tenth)
Use the function to predict the life expectancy of males in 2008. E(17)=___
(round to the nearest tenth)

Answer by stanbon(75887)

In 1991, the life expectancy of males in a certain country was 66.3 years. In 1995, it was 68.9 years. Let E represent the life expectancy in year t and let t represent the number of years since 1991.
The linear function E(t) that fits the data is E(t)=___t + ____ (round to the nearest tenth)
Use the function to predict the life expectancy of males in 2008. E(17)=___
---------------------
You have two points relating year and life expectancy: (0,66.3) and (4,68.9)
slope: (68.9-66.3)/4 = 0.65
intercept: E(0) = 66.3
-------------------------------
Equation: E(t)=0.65t+66.3
---
E(17) = 0.65*17+66.3 = 77.35
====================================
Cheers,
Stan H.